# A Gated Residual Kolmogorov-Arnold Networks for Mixtures of Experts

### v1

@@ I Introduction
Introduced by [1, 2], the Mixture of Experts (MoE) is one of the most popular machine learning techniques. It involves the combination of multiple neural networks, where each network called “expert”, focusing on specific parts of the input. A gating network is used to assign different weights to these "experts". This gating mechanism is deciding which expert should be most trusted for a given input. This approach helps in improving the overall performance of the system by leveraging the strengths of individual neural networks specialized in different tasks. The MoE has since been applied to various domains such natural language processing [3, 4], computer vision [5, 6, 7], and reinforcement learning [8]. Time series data can exhibit complex, non-linear patterns and regime shifts [9] . The MoE’s ability to combine multiple specialized expert networks can help capture these complex temporal dynamics observed in financial time series [10]. Despite the increase in the availability of dataset for time series forecasting, it reamains a very complex task. Powerful models have been proposed through the years [11, 12]. Some research proposed dynamic system to model different existing states within a time series [13, 14, 9]. In this paper, we explore the Mixture of Experts (MoE) using a novel neural networks architecture embedded in a gating mechanism. This new gating mechanism is built over Kolmogorov-Arnold Networks [15], allowing to estimate more accurately weights for each neural networks denoted $f_{i,\theta}\text{ and }i\in[0;n]$.
Recently, Liu et al. [15] released the Kolmogorov-Arnold Netowrks (KANs) a promising architecture, introduced as an alternative for MLPs. To dive into the architecture of the model, we redirect the reader to [15]. In our previous work we have adapted KANs for time series forecasting [16], other researchers also proposed an adaptation for time series [17]. According to [18] KANs outperforms conventional MLPs in real-world forecasting task. In a previous work we also proposed a temporal transformer architecture using TKANs [19]. Other extension using wavelets [20] have been proposed in the meantime. TKANs aim to develop a framework with two key functionalities: handling sequential data and managing memory. To do so, in our previous work [16] we enriched the initial model adding a recurring layer of KANs to introduce RKANs. Temporal Kolmogorov-Arnold Networks (TKANs) is an upgraded version of LSTM [21] using Kolmogorov-Arnold Networks (KANs). They rely on RKANs for the management of short term memory as well as a cell state; for further details, we refer the reader to [16].
@@ II-C Experts
We denote $f_{k,\theta_{k}}(.)$ an highly parametrized function which describe the sucession of operations of a neural networks such GRU, LSTM and TKAN. Let us denote $m$ the total number of experts, $k\in\{1,...,m\}$. The role of the expert in the framework is the learn from pattern within the sequence in the dataset. Each ouput of the experted is denoted $\tilde{y}_{k}$ and obtained simutaneously from the transformed input defined Eq. (1)
@@
All the estimated about per experts will be stacked in $\tilde{y}_{i}$
@@ II-D Output
The ouput of this transformation will feed the GRKAN layer to obtained the weight These ouputs will be converted to weights,
@@ III Learning
The first learning task is inspired by the one in [16], [19] and [22] with a few modifications. The task is about predicting market trading notional on multiple steps ahead on one asset using the past values of the asset but also others. We chose this task as it’s a challenging one to perform due to the very noisy series but also the internal patterns in the traded volume like strong auto-correlation and seasonality.
@@
The model are kept simple to study the effect of MoE and KAMoE on them. As we are on a sequential prediction tasks we used the two most standard RNN layer that are Gated Recurrent Unit (GRU), and Long Short Term Memory (LSTM). Each model is done by applying a first layer that return the full sequence, followed by a second that only returns the last element. Finally, this output is fully connected with a linear layer that has the same number of units that their is step to predict. The two hidden RNN layers are both composed of 100 units. Finally, to test the MoE and KAMoE effect we only apply it on the first hidden RNN layer, and is composed of 3 experts and used a sigmoid activation on the GRKAN outputs.
@@
For this task we compare the KAMoE effect on non-sequential model, we thus used very simple model which are an MLP and a KAN model. The MLP is composed of successive layers with ReLU activation, while not applying other activation to the KAN. We wanted to take the opportunity of this task to see the effect of the number of layers and hidden size for both, as well as being able to analyze wether the addition of the mixture framework do not increase results just by increasing the number of units. Finally we also compared the results with the one obtained with standard machine learning model using the one from scikit-learn as well as XGBoost.

### v2

@@ I Introduction
Introduced by [1, 2], the Mixture of Experts (MoE) is one of the most popular machine learning techniques. It involves the combination of multiple neural networks, where each network called “expert”, focusing on specific parts of the input. A gating network is used to assign different weights to these "experts". This gating mechanism is deciding which expert should be most trusted for a given input. This approach helps in improving the overall performance of the system by leveraging the strengths of individual neural networks specialized in different tasks. The MoE has since been applied to various domains such natural language processing [3, 4], computer vision [5, 6, 7], and reinforcement learning [8]. Time series data can exhibit complex, non-linear patterns and regime shifts [9] . The MoE’s ability to combine multiple specialized expert networks can help capture these complex temporal dynamics observed in financial time series [10]. Despite the increase in the availability of datasets for time series forecasting, it remains a very complex task. Powerful models have been proposed through the years [11, 12]. Some research proposed dynamic systems to model different existing states within a time series [13, 14, 9]. In this paper, we explore the Mixture of Experts (MoE) using a novel neural networks architecture embedded in a gating mechanism. This new gating mechanism is built over Kolmogorov-Arnold Networks [15], allowing for a more accurate estimation of weights for each neural network denoted $f_{i,\theta}\text{ and }i\in[0;n]$.
Recently, Liu et al. [15] released the Kolmogorov-Arnold Networks (KANs) a promising architecture introduced as an alternative for MLPs. To dive into the architecture of the model, we redirect the reader to [15]. In our previous work we have adapted KANs for time series forecasting [16], other researchers also proposed an adaptation for time series [17]. According to [18], KANs outperform conventional MLPs in real-world forecasting tasks. In a previous work we also proposed a temporal transformer architecture using TKANs [19]. Other extensions using wavelets [20] have been proposed in the meantime. TKANs aim to develop a framework with two key functionalities: handling sequential data and managing memory. To do so, in our previous work [16] we enriched the initial model by adding a recurring layer of KANs to introduce RKANs. Temporal Kolmogorov-Arnold Networks (TKANs) is an upgraded version of LSTM [21] using Kolmogorov-Arnold Networks (KANs). They rely on RKANs for the management of short-term memory as well as a cell state; for further details, we refer the reader to [16].
@@ II-C Experts
We denote $f_{k,\theta_{k}}(.)$ a highly parametrized function which describes the succession of operations of neural networks such GRU, LSTM and TKAN. Let us denote $m$ the total number of experts, $k\in\{1,...,m\}$. The role of the expert in the framework is the learn from patterns within the sequence in the dataset. Each output of the expert is denoted $\tilde{y}_{k}$ and obtained simultaneously from the transformed input defined Eq. (1)
@@
All the estimates per experts will be stacked in $\tilde{y}_{i}$
@@ II-D Output
The output of this transformation will feed the GRKAN layer to obtain the weight. These outputs will be converted to weights,
@@ III Learning
The first learning task is inspired by the one in [16], [19] and [22] with a few modifications. The task is about predicting market trading notional multiple steps ahead on one asset using the past values of the asset but also others. We chose this task as it’s a challenging one to perform due to the very noisy series but also the internal patterns in the traded volume like strong auto-correlation and seasonality.
@@
The model is kept simple to study the effect of MoE and KAMoE on them. As we are on a sequential prediction task, we used the two most standard RNN layers that are Gated Recurrent Unit (GRU), and Long Short Term Memory (LSTM). Each model is done by applying a first layer that returns the full sequence, followed by a second that only returns the last element. Finally, this output is fully connected with a linear layer that has the same number of units as there are steps to predict. The two hidden RNN layers are both composed of 100 units. Finally, to test the MoE and KAMoE effect, we only apply it on the first hidden RNN layer, which is composed of 3 experts and uses a sigmoid activation on the GRKAN outputs.
@@
For this task we compare the KAMoE effect on non-sequential model, we thus used very simple model which are an MLP and a KAN model. The MLP is composed of successive layers with ReLU activation, while not applying other activation to the KAN. We wanted to take the opportunity of this task to see the effect of the number of layers and hidden size for both, as well as being able to analyze whether the addition of the mixture framework does not increase results just by increasing the number of units. Finally we also compared the results with the one obtained with standard machine learning model using the one from scikit-learn as well as XGBoost.